OpenStudy (anonymous):

6x^6-48 factor completely

OpenStudy (anonymous):

im sorry what was that

OpenStudy (campbell_st):

factor out 6 and you are left with \[6(x^6 - 8)\] make a substitution \[u = x^2 \] and you have \[6(u^3 - 8)\] this is the difference of 2 cubes and can be factored hope it helps

OpenStudy (anonymous):

is it not factor bc it would be -6 + x6 =48

OpenStudy (campbell_st):

just read the solution I posted

OpenStudy (anonymous):

6(u^3)-4 i do not know if this is right

OpenStudy (campbell_st):

read the solution and all you need to do is factor the difference of 2 cubes

OpenStudy (anonymous):

i do not know how

OpenStudy (mathstudent55):

\(a^3 - b^3 = (a - b)(a^2 + ab + b^2)\)

OpenStudy (campbell_st):

6 multiplied by -8 = -48 that seems to be a mistake you have made

OpenStudy (anonymous):

im sorry

OpenStudy (anonymous):

6 (u)-8

OpenStudy (anonymous):

can you show me step by step

OpenStudy (anonymous):

?

OpenStudy (jhannybean):

What @campbell_st posted was if you make a substitution for x\(^6\) as u = x\(^2\), then you can think of your factored function as: \[6(\color{red}{(x^2)}^3-2^3)\] which is the difference of cubes. Difference of cubes: \(a^3 - b^3 = (a - b)(a^2 + ab + b^2)\) All you have to do is define \(a=x^2\) and \(b=2\) and fill in the form

OpenStudy (mathstudent55):

You need to use the pattern I wrote above for factoring a difference of cubes.

OpenStudy (jhannybean):

Forgot to tag @mathstudent55 for the nicely written formula posted above ^^

OpenStudy (jhannybean):

Forgot to tag @mathstudent55 for the nicely written formula posted above ^^

OpenStudy (mathstudent55):

@anthony6577 If you will pay attention, I will help you from beginning to end with this problem.

OpenStudy (anonymous):

im trying

OpenStudy (mathstudent55):

Ok, here we go. Factor: \(\large 6x^6 - 48\) The first step in factoring is always to try to factor a common factor.

OpenStudy (anonymous):

alright

OpenStudy (mathstudent55):

Here, there is a common factor in both terms. that common factor is 6, so we factor it out. \(\large = 6(x^6 - 8)\)

OpenStudy (mathstudent55):

The 6 will remain as a factor to the end, but we don;t need to worry about it. We need to figure out what to do with x^6 - 8.

OpenStudy (anonymous):

so the gcf is 6

OpenStudy (mathstudent55):

To factor a binomial, there are 3 possibilities: 1. A difference of squares. 2. A sum of cubes. 3. A difference of cubes.

OpenStudy (mathstudent55):

We look carefully at x^6 - 8 to figure out which of the three possibilities above, if any, it is.

OpenStudy (anonymous):

diffrent of cubs

OpenStudy (mathstudent55):

We notice that x^6 can be broken up in two ways: \(\large x^6 = (x^3)^2\) in this case, x^6 is the square of x^3, and is a square. \(\large x^2 = (x^2)^3 \) in this case, x^6 is the cube of x^2 and is a cube.

OpenStudy (mathstudent55):

We don't know yet whether we are dealing with cubes or squares. Now we look at 8.

OpenStudy (mathstudent55):

With 8, we notice that \(\large 8 = 2^3\) 8 is indeed a cube, it is the cube of 2.

OpenStudy (mathstudent55):

Since 8 is a cube, we need to treat x^6 as a cube also, so we have a difference of cubes.

OpenStudy (mathstudent55):

Now we know we are dealing with a difference of cubes, so we write x^6 - 8 as a difference of cubes clearly showing the cubes.

OpenStudy (anonymous):

so the anser would be 8+2^3

OpenStudy (mathstudent55):

\(\large =6((x^2)^3 - 2^3) \)

OpenStudy (mathstudent55):

No. We're not there yet. We now need to use the pattern of factoring a difference of cubes I wrote above.

OpenStudy (mathstudent55):

\(\large a^3 - b^3 = (a - b)(a^2 + ab + b^2)\)

OpenStudy (anonymous):

this pint threw me off

OpenStudy (mathstudent55):

The way a difference of cubes is factored is to first find what the cubes are cubes of. x^6 is the cube of x^2 8 is the cube of 2

OpenStudy (anonymous):

so 2

OpenStudy (mathstudent55):

Now you need two sets of parentheses. In the first one, you need the difference of the quantities without cubes. (x^2 - 2)

OpenStudy (anonymous):

alright

OpenStudy (mathstudent55):

In the second set of parentheses, you need the square of the first one: (x^2)^2 = x^4 Then add to it the product of the two0: x^2 * 2 = + 2x^2 Then add to that the square of the second one: 2^2 = + 4 Now fill in the second set of parentheses with these 3 quantities: (x^4 + 2x^2 + 4)

OpenStudy (mathstudent55):

Finally, we can write the complete solution: \(\large 6x^6 - 48\) \(\large =6(x^6 - 8)\) \(\large = 6((x^2)^3 - 2^3)\) \(\large = 6(x^2 - 2)(x^4 + 2x^2 + 4) \)

OpenStudy (anonymous):

so that is the work

OpenStudy (mathstudent55):

yes, and that is also the final answer

OpenStudy (anonymous):

that alot just to find out

OpenStudy (mathstudent55):

The solution is quick, but with all the explanations, it becomes long.

OpenStudy (anonymous):

thank you for spending the time to show me

OpenStudy (mathstudent55):

You are welcome.